40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 16:56:38 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades in progress 
    Reply  Wed Oct 9 17:27:35 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades  
       Reply  Wed Oct 9 22:05:55 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Fri Oct 11 17:37:08 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
             Reply  Sun Oct 13 13:29:09 2013, rana, Update, LSC, New lockin / sensing matrix screens 
                Reply  Mon Oct 14 17:40:15 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
                   Reply  Thu Oct 17 13:40:55 2013, Jenne, Update, LSC, New lockin / sensing matrix frequencies All_REFL_PRMI_17Oct2013.pdf
             Reply  Mon Oct 14 21:12:35 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Wed Oct 16 03:04:37 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
             Reply  Wed Oct 23 19:28:22 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
Message ID: 9225     Entry time: Wed Oct 9 17:27:35 2013     In reply to: 9222     Reply to this: 9227
Author: Jenne 
Type: Update 
Category: LSC 
Subject: LSC model sensing matrix upgrades  

The modifications to the LSC model are now complete, and the new model has been compiled, installed, and is running. The sensing matrix lockins are all in the c1cal model.  Masayuki is locking right now, so so far, things appear to be back to normal.

The longer version of the story, with all the detours and hiccups:

After several tries of deleting the GoTo and From flags / tags in the lockin area of the LSC model, and continually getting the "something is not connected" errors, I gave up and just drew several long lines.  So, in the new Sensing Matrix block (which is actually in c1cal, not c1lsc, but that's a story for the next paragraph), we should eventually make things back to the more clean flags situation, but for right now, it's working, with lots of lines everywhere.  I've tried to be very organized and clear about what lines go where, so that it's easy to see what's going on.

I eventually was able to compile c1lsc, and then installed it, and restarted the model.  Adding in 5x the lockin modules (we had 27, but now we have 5x27, so that we can look at the sensing matrix elements for every degree of freedom, and every photodiode, all at the same time) was too much for the poor lsc cpu.  I was consistently getting over 70usec per cycle, and was hitting a max of 77usec for the lsc cpu.  Both of those numbers are greater than the allowed 60usec.  So, I made the decision to put the whole sensing matrix / lockin stuff into the calibration model.  This means that I have 27+5 more IPC signals than I used to, but so far things seem to be okay (no rigorous testing yet).  (27 to send the 27 PD inputs over to the cal model, and another 5 to send the oscillator "clocks" from the cal model to the lsc model.)  The lsc model is now running faster than before (because there were 27 lockin modules in the model), at 24-28 usec, and the cal model is running at 39usec.

All seemed well and good, both the lsc and cal models compiled, but the lsc burt restore wasn't working.  Restarting the model did not successfully do a burt restore, and when I tried several different .snap files from today, and other times this month using burtgooey, I kept getting "NOT OK", and numbers weren't being restored into the epics channels.  A very few settings were restored, but most were not.  I ended up copying a .snap file from a few hours ago into a separate directory, then went in and by-hand removed all the lines referring to now non-existant lockin channels.  Burtgooey still told me "NOT OK", but settings seem to be restored, so I think it's okay.  I have not confirmed each and every one of the 10,000+ channels to ensure that the number is the same as the one in the .snap file, but as I glance around in the LSC screen and its dependants, all the numbers and buttons look about right.  

After all this stuff, Masayuki is locking both arms simultaneously in IR, as he prepares to test some new ALS scripts, so things seem okay for now.

ELOG V3.1.3-